A car's engine is turning the crankshaft at 5900 rev/min

rev/min .


what is a rotational speed?


1
Expert's answer
2021-05-30T13:32:14-0400

"ω = \frac{2π\times 5900}{60} = \frac{37052}{60}=617.53 \;rad/s"